OpenStudy (anonymous):

Jack usually mows his lawn in 6 hours. Marilyn can mow the same yard in 4 hours. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

2 hrs

OpenStudy (amistre64):

the idea is to determine how much work each one does in 1 "hour" as this case is

OpenStudy (amistre64):

J can do it all in 6 hours; how much can he do in 1 hour? 1/6 of it right?

OpenStudy (amistre64):

M can do it all in 4 hours, which means that in 1 hour she does 1/4 of it

OpenStudy (amistre64):

together in an hour they can do: 1/6 + 1/4 = (6+4)/24 = 10/24

OpenStudy (amistre64):

this tells us we divvy the yard up into 24 pieces and see how long it takes them to finish it all 10 * t = 24 t = 24/10 = 2.4 hours

OpenStudy (amistre64):

if I was british, I could convert that into minutes